### v1

# TRIM: Token-wise Attention-Derived Saliency for Data-Efficient Instruction Tuning

Abstract
Instruction tuning is essential for aligning large language models (LLMs) to downstream tasks and commonly relies on large, diverse corpora. However, small, high-quality subsets, known as coresets, can deliver comparable or superior results, though curating them remains challenging. Existing methods often rely on coarse, sample-level signals like gradients, an approach that is computationally expensive and overlooks fine-grained features. To address this, we introduce TRIM (Token Relevance via Interpretable Multi-layer Attention), a forward-only, token-centric framework. Instead of using gradients, TRIM operates by matching underlying representational patterns identified via attention-based "fingerprints" from a handful of target samples. Such an approach makes TRIM highly efficient and uniquely sensitive to the structural features that define a task. Coresets selected by our method consistently outperform state-of-the-art baselines by up to 9% on downstream tasks and even surpass the performance of full-data fine-tuning in some settings. By avoiding expensive backward passes, TRIM achieves this at a fraction of the computational cost. These findings establish TRIM as a scalable and efficient alternative for building high-quality instruction-tuning datasets.
TRIM: Token-wise Attention-Derived Saliency for Data-Efficient Instruction Tuning
Manish Nagaraj, Sakshi Choudhary, Utkarsh Saxena, Deepak Ravikumar, Kaushik Roy Department of Electrical and Computer Engineering, Purdue University {mnagara, choudh23, saxenau, dravikum, kaushik}@purdue.edu
1 Introduction
Large language models (LLMs) have become the de facto standard for a wide range of tasks, but their full potential is unlocked only after a critical, computationally expensive step: instruction tuning. While large, diverse instruction-tuning corpora have shown to be effective (Ouyang et al., 2022; Chung et al., 2024), a fundamental challenge persists: not all data is created equal. Recent findings have demonstrated that carefully curating a small, high-quality subset, or a “coreset”, can not only match but even surpass the performance of finetuning on the full dataset, all while using a fraction of the compute (Xia et al., 2024; Zhang et al., 2025a).
Current solutions often use influence-based methods to estimate sample-level data importance, relying on either gradients (Xia et al., 2024; Zhang et al., 2025b, c) or Hessians (Kwon et al., 2024; San Joaquin et al., 2024; Lin et al., 2024a). While effective, these are prohibitively expensive for large-scale data selection, requiring multiple backward passes, extensive memory, or a batch size of one to obtain per-sample gradients Pruthi et al. (2020). Representation similarity metrics (Hanawa et al., 2021; Zhang et al., 2018) alleviate the computational burden but still operate at the sample level. This coarse granularity introduces two key issues. First, a length bias, where deriving sample-wise scores from aggregated token-level metrics (e.g., loss, gradient magnitudes) creates a dependency on sequence length. Moreover, recent work shows that gradient norms in instruction-tuning samples are negatively correlated with sequence length. Xia et al. (2024). Second, a loss-centric bias, the next-token prediction objective distributes credit and blame uniformly across all tokens. In practice, however, softmax attention is inherently sparse, with performance often determined by a small subset of influential tokens Zhang et al. (2023); Ge et al. (2023); Lin et al. (2024b). Token-agnostic importance measures based solely on loss or gradients can misrank samples whose key signals are sparse and highly localized at a few tokens.
In principle, the most faithful measure of influence would come from token-wise gradients (or Hessians), but this is computationally infeasible in practice. We therefore aim to develop efficient proxies that preserve token-level resolution while avoiding the prohibitive cost of gradient computation. This leads us to a central question: “Given only a few target samples for a downstream task, how do we efficiently identify high-impact instruction data for task-specific fine-tuning?"
Figure 1: Overview of TRIM. (a) From multi-layer attention, we derive an aggregated token saliency signal that combines row (allocation sharpness) and column (received attention) signals. (b) Using a small target set, we compute one fingerprint per token class by saliency-weighted averaging of last-layer hidden states, capturing task-defining patterns (Section˜3.1). (c) For each candidate sample, token states are matched to their class fingerprints (cosine similarity), and token scores are pooled into a single relevance score for ranking (Section˜3.2); top-ranked samples form the instruction-tuning coreset.
We propose a shift from the traditional sample-level, loss-centric paradigm to a token-level, representation-centric one, where we leverage the model’s hidden states, rich numerical vectors that capture a token’s meaning in context. Our core insight is that a sample’s true utility often resides not in the entire sequence but in a few informative tokens. To that effect, we introduce TRIM (Token Relevance via Interpretable Multi-layer Attention), a novel method that identifies influential data by matching the underlying attention patterns of a target task.
As shown in Figure˜1, TRIM is a two-stage pipeline requiring only forward passes. Stage I constructs lightweight token fingerprints that capture the essence of a target task from a small set of samples. We begin by computing an aggregated token saliency score that combines two complementary signals: row saliency, which measures how sharply a token distributes its own attention, and column saliency, which measures how much attention it receives from other tokens (panel a). For each token class, we then compute a fingerprint as the saliency-weighted mean of the last-layer hidden states of all target-set tokens in that class (panel b), yielding a concise summary of what that class looks like to the model. Stage II uses these fingerprints to score samples in the large candidate corpus. For each sample, we compare each token’s last-layer hidden state to its class fingerprint via cosine similarity (panel c), then pool the resulting token-level scores into a single relevance score that reflects the sample’s alignment with the target task. Samples with the highest scores are selected to form the final coreset, comprising training examples whose token-level representations most strongly align with the target task’s structural patterns.
This design yields three advantages. First, TRIM delivers superior performance and efficiency: with only 5–10 samples from target task, it surpasses state-of-the-art methods (Xia et al., 2024; Zhang et al., 2025b; Yang et al., 2024) while running orders of magnitude faster, owing to its forward-only design that avoids per-sample gradient computation (Sections˜4.1 and 5). Second, it offers high structural fidelity by matching core representational patterns (e.g., syntax, mathematical operators), which benefits challenging downstream tasks (Section˜4.2). Finally, its token-level scoring inherently mitigates length bias, a common failure mode of sample-level methods (Section˜4.4).
Our contributions are:
  • A Token-Centric Framework. We introduce TRIM, to our knowledge the first token-level, representation-centric data-selection method that leverages contextualized token representations via attention-derived aggregated saliency.
  • A Scalable and Efficient Algorithm. TRIM is a forward-only algorithm that is orders of magnitude faster than gradient-based alternatives, enabling practical and efficient data selection over massive instruction corpora.
  • State-of-the-art empirical performance. With only a 5% coreset and 5-10 target samples, TRIM outperforms state-of-the-art methods by up to 9% on downstream reasoning tasks, and even exceeds full-data fine-tuning on some tasks.
2 Related Work
The challenge of curating coresets for instruction tuning has motivated a wide body of work on automated data selection (Cao et al., 2024; Chen et al., 2024; Zhou et al., 2023; Ding et al., 2023). Existing methods can be grouped into several paradigms. Influence-Based Selection methods estimate data importance by quantifying the causal effect of a candidate example on a target task. As exact influence functions (Koh and Liang, 2017) are infeasible for LLMs, recent works use scalable approximations. Some methods leverage Hessians Kwon et al. (2024); San Joaquin et al. (2024); Lin et al. (2024a), while other first-order alternatives track gradient similarity across checkpoints, as exemplified by the state-of-the-art LESS method (Xia et al., 2024). Forward-only approaches like CLD (Nagaraj et al., 2025) measure influence by correlating per-sample loss trajectories with validation set dynamics, though at the cost of reduced selection quality compared to gradient-based methods. Training-Dynamics-Based Selection methods identify informative examples without an explicit target set. Approaches like TAGCOS (Zhang et al., 2025b) and STAFF (Zhang et al., 2025c) use gradient features from warmup checkpoints or surrogate models. Other efficient, forward-only methods such as S2L use small-model losses to guide selection for larger models (Yang et al., 2024). In contrast, TRIM uses targeted, context-aware token-level selection from a few task samples; outperforming the task-agnostic methods (Table˜1).
Token-Centric Selection. Recognizing the limitations of sample-level metrics, the most recent work has shifted towards token-level signals. This emerging paradigm addresses goals such as improving long-context recall (Chen et al., 2025) or performing general-purpose dataset filtering through complex, multi-stage schemes (Fu et al., 2025). TRIM contributes to this emerging token-centric paradigm but is distinct in its goal and mechanism. Unlike methods for general pruning or long-context analysis, TRIM is specifically built for efficient, targeted selection guided by a small set of samples. It utilizes a novel lightweight saliency heuristic derived from multi-layer attention flow to quantify token importance (Abnar and Zuidema, 2020). By adapting this principle into a single-pass algorithm, TRIM offers a uniquely scalable solution, while outperforming state-of-the-art coreset generation techniques. A more comprehensive review of related works is provided in appendix˜A.
3 Methodology
Setup and Notation.
We begin with a pretrained language model $\mathbf{M}_{0}$. Our goal is to select a compact coreset $\mathcal{C}$ from a large and diverse instruction-tuning corpus $\mathcal{S}=\{s_{1},\dots,s_{|\mathcal{S}|}\}$, which serves as the candidate pool for fine-tuning $\mathbf{M}_{0}$ on a target domain $\mathcal{T}$. Coreset selection is guided by a small validation set $\mathcal{T}_{\text{val}}$, consisting of as few as 551010 samples from $\mathcal{T}$. Similar to prior works (Xia et al., 2024), we perform a brief warmup phase to adapt $\mathbf{M}_{0}$ to the source distribution and stabilize the attention readouts and token representations used for sample scoring. Specifically, we fine-tune $\mathbf{M}_{0}$ on a small random subset $\mathcal{S}_{\text{warmup}}\subset\mathcal{S}$ (about 5%), which produces the scoring model $\mathbf{M}_{\text{warmup}}$. All subsequent computations, i.e., attention-based saliency, fingerprint construction from $\mathcal{T}_{\text{val}}$, and candidate scoring over $\mathcal{S}$, are performed using only forward passes through $\mathbf{M}_{\text{warmup}}$. TRIM then executes a forward-only, two-stage pipeline guided by $\mathcal{T}_{\text{val}}$. In Stage I (Section˜3.1), it reads attention signals from $\mathbf{M}_{\text{warmup}}$ to construct token-wise, task-defining fingerprints from the examples in $\mathcal{T}_{\text{val}}$. In Stage II (Section˜3.2), each candidate $c\in\mathcal{S}$ is scored by comparing its token representations to the fingerprints, aggregating these comparisons into an example-level score, and selecting the top-ranked examples to form the final coreset $\mathcal{C}$.
3.1 Stage I: Building Saliency-Weighted Token Fingerprints
Given the target validation set $\mathcal{T}_{\text{val}}$, TRIM constructs a dictionary of token fingerprints. This is achieved through forward passes over $\mathbf{M}_{\text{warmup}}$, where we extract attention signals from the final $L$ layers.
Attention-Derived Token Saliency.
We define a token’s saliency by integrating complementary signals from the row (query) and column (key) dimensions of the attention feature map (Figure˜1a).
Row Saliency. We score a token at position $i$ by the sharpness of its attention distribution. Tokens that focus strongly on a few positions (keys) are considered more salient, while those that spread attention broadly are less informative. We capture this notion of attention sharpness as a proxy for token importance using the entropy of the attention distribution. For layer $l$, head $h$ and sequence length $T$, let $A^{l,h}\in\mathbb{R}^{T\times T}$ denote the attention matrix. We drop $h$ and $l$ notation for simplicity, and compute the entropy for $i^{th}$ row (query) as follows:
$H_{i}\;=\;-\sum_{j}A_{i,j}\,\log\!\big(A_{i,j}+\varepsilon\big).$ (1)
Here, $\varepsilon$ is added for numerical stability, and we further normalize this entropy by the number of non-masked keys to obtain the row saliency score:
$q_{i}\;=\;1\;-\;\frac{H_{i}}{\log\,\big|\{j:\,A_{i,j}>0\}\big|},$ (2)
This is aggregated across the last $L$ layers and $H$ heads to produce a final row saliency score:
$Q_{i}\;=\;\frac{1}{LH}\sum_{l=1}^{L}\sum_{h=1}^{H}q^{l,h}_{i}\;\in[0,1].$ (3)
Column Saliency. The column saliency score is computed by aggregating all attention weights for each column (key) token. This signifies which tokens are attended strongly by others and are hence more crucial for final performance. We quantify this by measuring average attention weights for $j^{th}$ column as follows:
$k_{j}\;=\;\frac{\sum_{i}A_{i,j}}{\big|\{i:\,A_{i,j}>0\}\big|}$ (4)
Similar to row saliency, we aggregate this across heads and layers:
$K^{\text{raw}}_{j}\;=\;\frac{1}{LH}\sum_{l=1}^{L}\sum_{h=1}^{H}k^{l,h}_{j}$ (5)
Min-max scaling is applied to normalize $K^{\text{raw}}_{j}$:
$K_{j}\;=\;\frac{K^{\text{raw}}_{j}-\min K_{j}^{\text{raw}}}{\max K_{j}^{\text{raw}}-\min K_{j}^{\text{raw}}+\varepsilon}\;\in[0,1].$ (6)
Here, $\varepsilon$ ensures numerical stability.
Aggregated Token Saliency. We perform a weighted averaging for the row and column saliency scores to obtain the aggregated saliency for the $i^{th}$ token:
$\alpha_{i}\;=\;w_{Q}\,Q_{i}\;+\;w_{K}\,K_{i}.$ (7)
This convex combination balances the complementary roles of row and column saliency, with higher $\alpha_{i}$ indicating greater token importance. For simplicity, we choose equal weights ($w_{Q}=w_{K}=0.5$) for our experiments.
Fingerprint Construction.
For each sample $v\in\mathcal{T}_{\text{val}}$ in the target validation set, we construct token fingerprints using the aggregated token saliency scores. Fingerprints are defined at the token-class level, where each class corresponds to a unique token in the vocabulary Wu and Papyan (2024). Let $t$ denote a token class, and $h_{v,i}$ be the last layer’s hidden state for token $t$ at position $i$ in a sample $v$. We collect the set of all such occurrences:
$O_{t}\;=\;\{(v,i):\,v\in\mathcal{T}_{\text{val}},\ \text{class}(v_{i})=t\}.$ (8)
The fingerprint for class $t$, denoted $f_{t}$, is then defined as the saliency-weighted average of the normalized hidden states $\hat{h}_{v,i}=\frac{h_{v,i}}{\|h_{v,i}\|_{2}}$:
$f_{t}\;=\;\frac{\sum_{(v,i)\in O_{t}}\alpha_{v,i}\,\hat{h}_{v,i}}{\left\|\sum_{(v,i)\in O_{t}}\alpha_{v,i}\,\hat{h}_{v,i}\right\|_{2}}.$ (9)
The resulting dictionary $\mathcal{F}=\{f_{t}\}$ compactly summarizes the validation set at the token level, emphasizing informative tokens through their saliency scores and contextual hidden representations.
3.2 Stage II: Candidate Scoring and Selection
With the aid of the fingerprint dictionary $\mathcal{F}$, TRIM assigns a score to each candidate sample $c\in\mathcal{S}$. To do so, we obtain the last-layer hidden states $h_{c,j}$ via a forward pass through $\mathbf{M}_{\text{warmup}}$. Intuitively, a candidate is valuable if its token representations align with the salient, task-defining fingerprints derived from $\mathcal{T}_{\text{val}}$. For each token at position $j$ in candidate $c$, let $t_{j}$ be its token class. We then compute its similarity score $s_{j}$ by measuring the cosine similarity between its normalized hidden state, $\hat{h}_{c,j}=\tfrac{h_{c,j}}{\|h_{c,j}\|_{2}}$, and the corresponding class fingerprint $f_{t_{j}}$.
$s_{j}\;=\;\cos\!\big(\hat{h}_{c,j},\,f_{t_{j}}\big).$ (10)
Handling Non-fingerprinted Token Classes.
If a token class $t_{j}$ does not appear in $\mathcal{T}_{\text{val}}$, it will not have a corresponding fingerprint in $\mathcal{F}$. For these unseen classes, we find the nearest fingerprinted class in the input-embedding space via cosine similarity and apply a penalty factor $\lambda\in(0,1]$ to down-weight its contribution. Let $E$ denote the input-embedding matrix of $\mathbf{M}_{\text{warmup}}$ and $e_{t}$ be the unit-normalized embedding for class $t$. The nearest fingerprinted class is:
$\bar{t}_{j}\;=\;\arg\max_{t\,\in\,\mathcal{F}}\;\cos\!\big(e_{t_{j}},\,e_{t}\big),$ (11)
and the penalized similarity score is:
$s_{j}\;=\;\lambda\cdot\cos\!\big(\hat{h}_{c,j},\,f_{\bar{t}_{j}}\big).$ (12)
Robust Sample Score.
The final score for a candidate $c$ is aggregated over its token set $M(c)$ using a robust pooling function that combines the mean and maximum of its token scores:
$S(c)=w_{\mu}\tfrac{1}{|M(c)|}\!\sum_{j\in M(c)}\!s_{j}+w_{m}\max_{j\in M(c)}s_{j}$ (13)
The mean term captures overall informativeness across tokens, while the max term highlights the presence of highly crucial tokens. For simplicity, we fix $w_{\mu}=w_{m}=0.5$.
Coreset Selection.
Finally, we rank all candidate examples $c\in\mathcal{S}$ by their scores $S(c)$ and select the top-ranking examples to form the coreset $\mathcal{C}$. Since this scoring stage only requires a single forward pass per candidate, its runtime is linear in the corpus size once the fingerprints are constructed. For detailed pseudocode, please refer to Appendix˜B.
4 Experiments
We assess TRIM on three axes: (i) budgeted accuracy across general-domain multiple-choice reasoning benchmarks (Section˜4.1) and a low-overlap out-of-domain math benchmark (Section˜4.2); (ii) cross-model/scale transfer (Section˜4.3); and (iii) robustness to length bias (Section˜4.4).
Experimental Setup.
We select from a large candidate pool of 270k instruction-tuning examples (Dolly (Conover et al., 2023), CoT (Wei et al., 2022), Oasst1 (Köpf et al., 2023), Flan_v2 (Longpre et al., 2023)) to fine-tune models for general-domain multiple-choice reasoning benchmarks (CommonsenseQA (Talmor et al., 2019), SocialIQA (Sap et al., 2019), HellaSwag (Zellers et al., 2019)) and a mathematical reasoning benchmark (GSM8K (Cobbe et al., 2021)). We use Llama-3.2-1B for the main comparisons and Llama-2-7B (Touvron et al., 2023) for GSM8K. For transferability, we evaluate how coresets selected by a Llama-3.2-1B scorer are applied to larger or different architectures, specifically Llama-3.1-8B (Grattafiori et al., 2024) and Mistral-7B (v0.3) (Jiang et al., 2023). All models are trained with LoRA (Hu et al., 2022), and selection of candidate samples is guided by 5–10 target validation samples per task. The full setup and hyperparameter details are provided in Appendices DE, and F.
Baselines.
We group coreset methods by mechanism and defer the details to Appendix C. (i) Heuristics: Random, lexical retrieval (BM25 (Robertson et al., 2009)), and $N$-gram importance resampling (DSIR (Xie et al., 2023)). (ii) Forward-only training dynamics: CLD (correlation of loss differences) (Nagaraj et al., 2025) and S2L (loss trajectory clustering) (Yang et al., 2024). (iii) Representation-based: RDS (Hanawa et al., 2021). (iv) Gradient-based: LESS (Xia et al., 2024) and TAGCOS (Zhang et al., 2025b). All methods share a standardized warmup-and-selection pipeline (Appendix F). We warm up Llama-3.2-1B on a random 5% subset of the source dataset using LoRA. TRIM/TAGCOS score candidates using the final warmup checkpoint, LESS/RDS aggregate across checkpoints, and CLD/S2L use trajectories collected during full-dataset training. For context, we also report Pretrained and Full-data baselines.
4.1 Performance Comparison
We evaluate all methods at a fixed 5% budget. Each method selects its coreset from the candidate pool, which is then used to fine-tune a fresh Llama-3.2-1B model with LoRA, initialized from the pretrained checkpoint (i.e., not the warmed-up scorer).
Coreset Method CommonsenseQA SocialIQA HellaSwag Average Mean
[…]
Random 34.05 (1.2) 43.84 (0.1) 48.21 (0.1) 42.03
BM25 (Robertson et al., 2009) 38.88 (0.7) 44.41 (0.3) 48.76 (0.1) 44.02
DSIR (Xie et al., 2023) 37.16 (0.7) 44.38 (0.5) 48.75 (0.2) 43.43
S2L (Yang et al., 2024) 34.10 (0.1) 43.32 (0.2) 48.57 (0.2) 41.99
RDS (Hanawa et al., 2021) 36.16 (1.1) 43.77 (0.1) 48.43 (0.3) 42.79
CLD (Nagaraj et al., 2025) 33.10 (0.4) 43.25 (0.5) 48.35 (0.1) 41.56
LESS (Xia et al., 2024) 39.10 (0.7) 44.52 (0.3) 49.01 (0.1) 44.21
TAGCOS (Zhang et al., 2025b) 34.72 (0.8) 43.70 (0.2) 48.76 (0.1) 42.39
[…]
Table 1: Accuracy (%) for coreset methods on Llama-3.2-1B with a 5% coreset. Results are means over 3 seeds (standard deviations in parentheses). The final column is the macro-average across tasks.
Results and Takeaways.
As summarized in Table˜1, TRIM attains the highest mean accuracy of 45.37%, exceeding the next-best method, LESS (44.21%), by over one point. Notably, TRIM’s 5% coresets are the only ones that surpass full-data fine-tuning on two tasks, SocialIQA (46.26% vs. 45.04%) and HellaSwag (49.08% vs. 48.55%).
Figure 2: TRIM vs. top baselines on Llama-3.2-1B across coreset budgets for CommonsenseQA (left) and SocialIQA (right). TRIM maintains a consistent advantage and, on SocialIQA, exceeds the full-data baseline with coresets as small as 1%.
Beyond the 5% budget, fig.˜2 shows TRIM consistently outperforming LESS and TAGCOS across coreset sizes. On SocialIQA, it surpasses full-data at 1% and peaks at 2.5%; on CommonsenseQA, it exceeds full-data at larger budgets ($20\%$).
While heuristic lexical methods (BM25, DSIR) track surface overlap; training-dynamics (CLD, S2L), representation-based (RDS), and gradient-based (LESS, TAGCOS) remain sample-level and thus coarse. On the other hand, TRIM’s token-centric scoring targets critical tokens, enabling a more precise selection. We provide an ablation study over our hyperparameters in Table G.
4.2 Low-Overlap Out-Of-Domain: Mathematical Reasoning
We stress test TRIM’s core hypothesis: token-level representational matching can surface training data that imparts structural reasoning skills even without topical overlap. Concretely, we ask each method to select a coreset for GSM8K from our general-purpose candidate instruction-tuning pool (not curated for math). We then fine-tune Llama-2-7B on the selected data and evaluate on GSM8K. This setting is challenging; fine-tuning on the entire corpus yields only 30.25%, so success requires identifying examples with latent structures (e.g., stepwise logic, numerical manipulation) rather than superficial topical cues. We include three strong sample-wise baselines: LESS, S2L, and TAGCOS.
[…]
Random 15.43 (0.80) 18.21 (0.6)
LESS (Xia et al., 2024) 17.34 (1.2) 20.72 (0.4)
S2L (Yang et al., 2024) 16.48 (0.4) 18.55 (0.3)
TAGCOS (Zhang et al., 2025b) 17.23 (0.2) 18.72 (0.4)
[…]
Table 2: Out-of-domain data selection for GSM8k on Llama-2-7B. The table shows exact-match accuracy (%) for coresets of 1% and 5% selected from a non-math corpus. All scores are the mean of 3 seeds.
Results and Takeaways.
As shown in Table˜2, a 5% TRIM coreset reaches 29.52%, nearly matching full-data fine-tuning (30.25%) and outperforming the next best method (LESS) by $\sim$8.8%. Even at a 1% budget, TRIM achieves 22.33% accuracy, remaining well ahead of alternatives. This stems from strong structural fidelity: by matching token-level fingerprints, rather than sample-level losses or gradients, TRIM retrieves examples that convey a style of reasoning (e.g., enumerations, logical connectives, stepwise instructions). These signals act as transferable proxies for chain-of-thought structure in GSM8K, yielding a higher performance that sample-level metrics fail to capture.
4.3 Coreset Transferability
We assess whether a small scorer can curate data that transfers to larger models. Using Llama-3.2-1B as the scorer, we select a single 5% coreset and reuse it to fine-tune two larger target models: Llama-3.1-8B (same model family) and Mistral-7B (v0.3) (different model family). We compare (i) TRIM-Transfer, the coreset chosen by the 1B scorer, and (ii) TRIM-Oracle, a coreset selected using the target model itself as the scorer. The oracle serves as an in-model reference point.
Llama-3.1-8B Mistral-7B (v0.3)
Coreset Method CSQA SIQA HS Avg. CSQA SIQA HS Avg.
[…]
TRIM-Transfer (from 1B) 76.38 (0.9) 51.95 (0.5) 64.19 (0.1) 64.17 75.98 (0.9) 54.84 (0.8) 65.35 (0.2) 65.39
TRIM-Oracle (in-model) 75.40 (0.2) 51.11 (0.9) 63.93 (0.1) 63.48 75.03 (0.4) 54.12 (0.4) 65.34 (0.1) 64.83
Table 3: Cross-Model and Cross-Scale Transferability of Coresets Selected by a 1B Scorer. All scores are the mean Accuracy (%) across three seeds. Datasets are: CommonsenseQA (CSQA), SocialIQA (SIQA), and HellaSwag (HS).
Results and Takeaways.
As shown in Table˜3, coresets selected by the 1B scorer transfer with high fidelity: on Llama-3.1-8B the transferred coreset reaches 64.17% vs. the oracle’s 63.48%, and on Mistral-7B it again edges the oracle (65.39% vs. 64.83%). Thus, a single coreset selected once by a small model can match, or even surpass, a per-target oracle. This strong transfer suggests that TRIM’s token-level fingerprints capture model-agnostic structure that generalizes across scale and architecture, enabling a lightweight scorer to curate high-quality coresets for large targets without sacrificing accuracy.
4.4 Mitigating Length Bias
A common failure mode in coreset selection is a bias toward shorter sequences. Sample-level heuristics can be confounded by sequence length, e.g., influence scores derived from gradient magnitudes tend to favor short examples with concentrated signals over longer prompts that contain richer reasoning (Xia et al., 2024). We test whether TRIM exhibits this bias by analyzing the length distribution of coresets selected for GSM8K by TRIM, LESS, and TAGCOS.
Figure 3: Length distribution of selected coresets for GSM8K. The histogram shows the percentage of selected samples by length bucket. Sample-level methods (LESS, TAGCOS) skew short, whereas TRIM selects a broader distribution with a higher mean length.
Results and analysis.
As shown in Figure˜3, LESS and TAGCOS skew toward shorter samples (mean lengths 323 and 259 tokens). In contrast, TRIM selects a broader distribution with many longer, more complex examples; its mean length is 446 tokens, $\sim$38% longer than LESS and $\sim$72% longer than TAGCOS. Since sample-level scoring can conflate brevity with quality, TRIM’s token-centric approach is effectively length-agnostic: by elevating important tokens wherever they occur, it yields diverse, informative coresets.
4.5 Task-Adaptive Data Selection
We analyze the distribution of selected examples across GSM8K and CommonsenseQA, as shown in Figure 4. TAGCOS maintains nearly identical selection patterns for both tasks, demonstrating its inability to adapt to task-specific requirements. In contrast, TRIM exhibits strong task-adaptive behavior: for CommonsenseQA, it predominantly selects from Flan_v2, leveraging its diverse instruction-following examples well-suited for question answering, while for GSM8K, it shifts toward the CoT dataset, prioritizing step-by-step reasoning essential for mathematical problem solving. Notably, LESS struggles to identify the importance of CoT data for GSM8K, while still favoring Flan_v2. This demonstrates that effective data selection requires not just quality assessment, but also the ability to recognize task-specific data characteristics, a capability that TRIM successfully captures. For more results, refer to Appendix˜H.
Figure 4: Distribution of selected training examples across data subsets for GSM8K and CommonsenseQA.
5 Computational Cost of Data Selection
We analyze the asymptotic cost of the selection stage (model preparation + candidate scoring). Let $N$ be the number of candidate examples and $T$ the number of training epochs used for model preparation (“warmup”). Let $f$ denote the FLOPs of a single forward pass of the scoring model. Following convention, a backward pass costs $\approx 2f$, so a full training step (forward + backward) is $\approx 3f$. We denote by $\gamma\in(0,1]$ the fraction of the dataset used for warmup and by $C$ the number of checkpoints whose states/representations are used during scoring. For a fair comparison, we assume all methods use a scoring model with forward-pass cost $f$. The cost to process the (small) target validation set of size $Q$ is negligible since $N\gg Q$ (in our setup, N>270N\!>\!270k and $Q\in\{5,10\}$).
Coreset Method Total selection cost
S2L (Yang et al., 2024) $\mathcal{O}(3fNT\;+\;fNC)$
LESS (Xia et al., 2024) $\mathcal{O}(3f\,\gamma NT\;+\;3fNC)$
TAGCOS (Zhang et al., 2025b) $\mathcal{O}(3f\,\gamma NT\;+\;3fN)$
RDS (Hanawa et al., 2021) $\mathcal{O}(3f\,\gamma NT\;+\;fNC)$
[…]
Table 4: Asymptotic selection cost under a common scoring model (forward cost $f$). The first term is model preparation (warmup), the second is candidate scoring.
With a fixed scoring model size, the drivers are (i) whether backward passes are required during scoring and (ii) how many checkpoints $C$ are used. LESS is most expensive at scoring time: it uses gradient features at multiple checkpoints, incurring a backward pass for each candidate at each checkpoint, yielding a $\mathcal{O}(3fNC)$ term in addition to warmup $\mathcal{O}(3f\,\gamma NT)$. Forward-only methods that do not compute gradients at scoring time (e.g., S2L, RDS) avoid the $3f$ factor at scoring but still scale linearly in $C$ via $\mathcal{O}(fNC)$. Importantly, S2L (Yang et al., 2024) must train on all $N$ candidates to record per-sample loss trajectories (the signal it clusters), so its preparation term is $\mathcal{O}(3fNT)$ rather than $\mathcal{O}(3f\,\gamma NT)$. TAGCOS scores at a single checkpoint (C=1C{=}1) but requires per-candidate gradients once, giving $\mathcal{O}(3fN)$ after warmup. In contrast, TRIM scores each candidate with a single forward pass at one checkpoint (C=1C{=}1), producing $\mathcal{O}(fN)$ after warmup. Consequently, TRIM is asymptotically the most efficient among the compared methods for large $N$, and its benefits increase when baselines require multiple checkpoints ($C\gg 1$) or gradient computation during scoring.
6 Conclusion
We introduced TRIM, a token-centric framework for efficient coreset selection in instruction tuning. By shifting from coarse, sample-level signals to fine-grained, attention-derived token representations, TRIM addresses two persistent challenges in data selection: computational cost and length bias. Through saliency-weighted fingerprints constructed from a handful of target samples, TRIM identifies training data that matches the structural patterns defining a task, rather than relying on expensive gradient computation or surface-level similarity. Our experiments demonstrate that TRIM consistently outperforms state-of-the-art methods across several downstream tasks, achieving up to 9% improvements with only 5% coresets. Notably, TRIM’s token-level scoring enables it to surface structurally relevant data even in low-overlap settings, as evidenced by near-full-data performance on mathematical reasoning using a general-purpose corpus. The method’s forward-only design delivers these gains at a fraction of the computational cost of gradient-based alternatives, and coresets selected by small scorers transfer effectively to larger models across architectures. TRIM’s ability to capture task-defining structure through attention signals offers a scalable path forward, enabling practitioners to curate smaller, higher-quality datasets.
Limitations
TRIM’s effectiveness depends on the quality and representativeness of the target validation set. With only 5-10 examples guiding selection, noise or bias in these samples can propagate into the fingerprints, potentially degrading coreset quality. Future work could explore ensemble approaches that aggregate fingerprints across multiple validation sets to improve robustness. Additionally, TRIM requires a target set, making it not directly compatible for pretraining data curation. However, the core principle of matching data-model interactions can be potentially extended to pretraining through task-agnostic quality signals, for instance, using broad capability benchmarks as proxy validation sets.
Appendix A Extended Related Work
This section provides extended details on prior work in data selection for instruction tuning, expanding upon the summary in section˜2.
Target-Aware Influence-Based Selection
Influence-based methods aim to quantify the causal effect of a candidate training example on model performance with respect to a target task. A foundational approach, Influence Functions (Koh and Liang, 2017), approximates the impact of upweighting a training point via second-order Hessian computations. As these are infeasible for LLMs, scalable surrogates have been proposed. DataInf (Kwon et al., 2024) and In2Core (San Joaquin et al., 2024) exploit the low-rank structure of LoRA finetuning to derive closed-form influence approximations, enabling per-sample scoring. DealREC (Lin et al., 2024a) instead estimates Hessian-vector products on surrogate models and adjusts with gradient norms as proxies for learning difficulty.
First-order alternatives avoid Hessians by tracking gradient similarity across training. TracIn (Pruthi et al., 2020) accumulates dot products of gradients between training and target examples over checkpoints. LESS (Xia et al., 2024), designed for instruction tuning, stores compact gradient features with optimizer-aware weighting, yielding strong coresets that can rival full-data finetuning. BIDS (Dai et al., 2025) further balances capabilities across tasks by normalizing influence scores prior to selection.
Training-Dynamics-Based Selection
Other methods focus on learning dynamics without explicit target sets, seeking data that is broadly informative. TAGCOS (Zhang et al., 2025b) clusters gradient features from warmup checkpoints to capture representativeness. STAFF (Zhang et al., 2025c) uses gradient norms from smaller surrogate models as effort-based scores.
Classical approaches like Forgetting (Toneva et al., 2018), EL2N, and GraNd (Paul et al., 2021) track prediction changes, error norms, and gradient norms, respectively, during early training. Moderate (Xia et al., 2022) score samples by the distance of the sample features to the class median. $\mathbb{D}^{2}$-Pruning (Maharana et al., 2024) uses a graph-based framework that uses the feature density as a measure of the diversity of samples and the prediction variance as the difficulty of the sample. Forward-only methods mitigate the cost of training on all candidates: S2L (Yang et al., 2024) leverages small-model losses to guide large-model selection, and DUAL (Cho et al., 2025) combines uncertainty with difficulty to stabilize pruning.
Heuristic and Scorer-Based Selection
This family includes efficient, task-agnostic methods. BM25 (Robertson et al., 2009) and DSIR (Xie et al., 2023) select based on lexical overlap or $n$-gram distributions, but lack target awareness. Recent works employ lightweight scorers or LLMs-as-judges to filter or rank examples. Examples include filtering low-quality responses (Chen et al., 2024), ranking samples for diversity (Ge et al., 2024), or targeting difficult instructions (Li et al., 2024; Liu et al., 2024a). These methods are efficient but depend on proxy notions of data quality.
Optimization-Based Selection
A principled line of work casts data selection as optimization. TSDS (Liu et al., 2024b) minimizes distribution alignment loss between coresets and target distributions. GREATS (Wang et al., 2024) greedily selects data by gradient alignment with validation batches. PDS (Gu et al., 2025) frames the task as optimal control, linking selection scores to downstream impact. These methods are elegant but computationally heavy.
Token-Centric and Attention-Based Selection.
Recognizing the limitations of sample-level metrics, the most recent work has begun to shift towards token-level signals. These methods leverage the internal mechanisms of Transformers to find important data. For instance, LADM uses attention scores to measure long-range dependencies for long-context data selection (Chen et al., 2025). T-SHIRT proposes a hierarchical filtering scheme for general-purpose data pruning, rather than for the task-specific, targeted selection that we address (Fu et al., 2025). Its complex pipeline first uses K-Means clustering for a coarse, document-level selection, then refines this by scoring individual tokens via iterative input perturbations, making it computationally intensive.
Connection to Linguistic Collapse
Linguistic Collapse (Wu and Papyan, 2024) extends Neural Collapse (Papyan et al., 2020) theory to causal LMs, documenting that with increased scale and appropriate regularization, the final-layer token representations align toward class-like centroids, with equinorm/equiangular structure, and that this geometry correlates with generalization. This effect is characterized at (or near) the end of pretraining. Our setting focuses on instruction finetuning with relatively shallow training trajectories and task-specific distributions. The relation is therefore tangential but pertinent: (i) Shared space. TRIM scores examples via token-level hidden-state similarities, precisely where collapse-like geometry manifests; stronger centroidal structure from pretraining can sharpen TRIM’s cosine signals for rare, task-informative tokens. (ii) Local reshaping under finetuning. Instruction FT can selectively reshape hidden states toward the target format; aggregating scores across warmup checkpoints (LR-weighted) makes TRIM less sensitive to any late-stage drift. (iii) Data selection as geometry control. By emphasizing TF-IDF-weighted, task-specific tokens, TRIM may preferentially retain examples that reinforce useful token centroids for the downstream task rather than inducing indiscriminate collapse. A full causal link between collapse metrics during pretraining and FT-time selection effectiveness remains an interesting open question.
Positioning TRIM
TRIM contributes to the emerging token-centric paradigm but is differentiated by its specific focus on efficient, targeted data selection. Unlike methods for general pruning or long-context analysis, TRIM uses a lightweight, single-pass heuristic to find data relevant to a small set of target examples.
Our approach shares a philosophical connection with token sparsification methods designed for model efficiency, most notably in computer vision. For example, DynamicViT (Rao et al., 2021) introduces a lightweight prediction module to progressively prune uninformative image patch tokens at multiple layers of a Vision Transformer. This drastically reduces FLOPs with minimal accuracy loss by focusing computation on the most salient parts of an input. Both TRIM and DynamicViT operate on the principle that focusing computation on a small subset of important tokens is a highly effective strategy for efficiency, one applies it to data selection, the other to the inference pass itself.
The specific mechanism of TRIM is grounded in prior work on attention analysis. The core idea of using raw attention scores as an explicit, guiding signal, rather than just for representation mixing, has proven successful in other NLP domains. For instance, the work on Self-Attention Guided Copy Mechanisms (Xu et al., 2020) used attention scores to inform the model which words from a source text were important enough to be copied directly into a summary, demonstrating their utility as a direct signal of importance. TRIM adapts this concept, using attention to signal which tokens are most important for defining a task-specific fingerprint.
Furthermore, our multi-layer approach to calculating salience is inspired by interpretability research. Work on "Quantifying Attention Flow" (Abnar and Zuidema, 2020) established the value of aggregating attention across all layers to build a complete information flow graph, thereby determining a token’s overall importance. While their graph-based max-flow algorithm is designed for deep interpretability, TRIM adapts the underlying principle into a fast heuristic (focus and centrality) suitable for a large-scale, practical data selection task.
Appendix B Pseudocode for TRIM
Algorithm 1 BuildFingerprints (Stage I)
1:Warmed model $\mathbf{M}_{\mathrm{warmup}}$, target val $\mathcal{T}_{\mathrm{val}}$, Number of layers $L$
2:Salience weights $(w_{Q},w_{K})=(\tfrac{1}{2},\tfrac{1}{2})$
3:Fingerprint dictionary $\mathcal{F}=\{f_{t}\}$
[…]
5:for each $v\in\mathcal{T}_{\mathrm{val}}$ do
6:  Run $\mathbf{M}_{\mathrm{warmup}}$ once (attn+states, last $L$)
7:  for each position $i$ in $v$ do
8:   Compute $Q_{i}$ and $\tilde{K}_{i}$ (eqs. in text)
9:   $\alpha_{i}\leftarrow w_{Q}Q_{i}+w_{K}\tilde{K}_{i}$
10:for each token type $t$ seen in $\mathcal{T}_{\mathrm{val}}$ do
11:  $O_{t}\leftarrow\{(v,i):\mathrm{type}(v_{i})=t\}$
12:  $\hat{h}_{v,i}\leftarrow h_{v,i}/\|h_{v,i}\|_{2}$
13:  $f_{t}\leftarrow\mathrm{normalize}\!\left(\sum_{(v,i)\in O_{t}}\alpha_{v,i}\hat{h}_{v,i}\right)$
14:  $\mathcal{F}[t]\leftarrow f_{t}$
15:return $\mathcal{F}$
Algorithm 2 ScoreCandidates (Stage II)
[…]
2:Pool weights $(w_{\mu},w_{m})=(\tfrac{1}{2},\tfrac{1}{2})$, Non fingerprinted penalty $\lambda\in(0,1)$
3:Scores $\{S(c)\}_{c\in\mathcal{S}}$
4:for each $c\in\mathcal{S}$ do
5:  Run $\mathbf{M}_{\mathrm{warmup}}$ once to get $\hat{h}_{c,j}$ for all $j$
6:  $M(c)\leftarrow\emptyset$
7:  for each position $j$ do
8:   $t_{j}\leftarrow\mathrm{type}(c_{j})$
9:   if $t_{j}\in\mathcal{F}$ then
10:     $s_{j}\leftarrow\cos(\hat{h}_{c,j},\,\mathcal{F}[t_{j}])$
11:   else
12:     $\bar{t}_{j}\leftarrow\arg\max_{t\in\mathcal{F}}\cos(e_{t_{j}},e_{t})$
13:     $s_{j}\leftarrow\lambda\cdot\cos(\hat{h}_{c,j},\,\mathcal{F}[\bar{t}_{j}])$
14:   $M(c)\leftarrow M(c)\cup\{j\}$
15:  if $|M(c)|=0$ then
16:   $S(c)\leftarrow$-$\infty$
17:  else
18:   $\mu\leftarrow\frac{1}{|M(c)|}\sum_{j\in M(c)}s_{j}$
19:   $m\leftarrow\max_{j\in M(c)}s_{j}$
20:   $S(c)\leftarrow w_{\mu}\mu+w_{m}m$
21:return $\{S(c)\}$
Algorithm 3 TRIM Framework
(Warmup $\rightarrow$ Stage I $\rightarrow$ Stage II $\rightarrow$ Selection)
1:Base LLM $\mathbf{M}_{0}$, pool $\mathcal{S}$, target val $\mathcal{T}_{\mathrm{val}}$, coreset size $K$
2:Coreset $\mathcal{C}\subset\mathcal{S}$ of size $K$
3:Warmup: sample $\mathcal{S}_{\mathrm{warmup}}\subset\mathcal{S}$ ($\sim$5%); fine-tune $\mathbf{M}_{0}$
[…]
8:return $\mathcal{C}$
We provide detailed pseudocode for the complete TRIM pipeline to complement the description in Section˜3.
The main driver, presented in Algorithm˜3, orchestrates the overall workflow: model warmup, fingerprint generation, candidate scoring, and final coreset selection. This algorithm relies on two core subroutines that correspond to the two stages of our method.
Stage I, implemented in BuildFingerprints (Algorithm˜1), details the logic from Section˜3.1. It processes the target validation set $\mathcal{T}_{\text{val}}$ using the warmed-up model $\mathbf{M}_{\text{warmup}}$. For each token, it computes the attention-derived salience $\alpha_{i}$ by combining the director ($Q_{i}$) and hub ($\tilde{K}_{i}$) scores. These salience values are then used to create the final, salience-weighted token fingerprints $\mathcal{F}=\{f_{t}\}$.
Stage II, detailed in ScoreCandidates (Algorithm˜2), executes the scoring process from Section˜3.2. For each candidate, it computes token-wise similarity to the fingerprints, applying a penalized backoff for out-of-vocabulary types. These token scores are then aggregated into a robust example score using a mean-max pooling strategy with a coverage bonus.
Together, these algorithms provide a concrete implementation of the forward-only TRIM pipeline, culminating in the selection of the coreset $\mathcal{C}$ for downstream finetuning.
B.1 Nuances in the Algorithm
Selective Token Scoring.
A key flexibility of TRIM is the ability to selectively fingerprint and score different parts of an example. This scope is a hyperparameter that can be set to all tokens, prompt-only, or response-only. For tasks where the crucial reasoning steps are in the generated answer (e.g., mathematical reasoning in GSM8K), we can restrict scoring to the response. Conversely, for tasks where the complexity lies in the prompt (e.g., Commonsense QA), we can focus only on the prompt. This allows TRIM to target the most informative part of the data for a given domain while scoring fewer tokens for more efficiency.
Defining the Scored Token Set.
If selective token scoring is done, then the token at position $j$ is included in $M(c)$ if and only if it falls within the chosen scoring scope (prompt, response, or all) and is not a special token (e.g., BOS, EOS, PAD). The token’s score $s_{j}$ is calculated as defined above, whether through a direct fingerprint match or OOV backoff.
$M(c)=\{\,j:\ \text{token }j\text{ is scored}\,\}.$ (14)
If, for a given candidate, this set is empty $\left(|M(c)|=0\right)$, we assign it a score of $S(c)=$-$\infty$. And the robust pooling operation will then include an additional penalty term called “coverage" ($\kappa(c)$).
$\mu(c)\;=\;$ $\frac{1}{|M(c)|}\sum_{j\in M(c)}s_{j},$
$m(c)\;=\;$ $\max_{j\in M(c)}s_{j},$
$\kappa(c)\;=\;$ $\frac{|M(c)|}{|c|}.$
$S(c)\;=\;w_{\mu}\,\mu(c)\;+\;w_{m}\,m(c)\;+\;\eta\,\kappa(c).$ (15)
The coverage term lightly rewards candidates where salient matches are not sparse and mitigates length bias. With a small weight $\eta\ll w_{\mu},w_{m}$, the coverage term acts as a tie-breaker, without suppressing genuinely strong single-token spikes.
Appendix C Baseline Data Selection Strategy
This section provides a detailed overview of the baseline methods used for comparison in our experiments.
Random
This is the simplest baseline, involving uniform random sampling of a subset of the desired size from the full candidate pool without replacement. It serves as a measure of the performance gain achieved by more sophisticated selection strategies.
BM25 (Robertson et al., 2009)
BM25 is a lexical retrieval method that scores candidate examples based on their textual similarity to the target validation set. It uses word frequency statistics, similar to TF-IDF (Sparck Jones, 1972), to rank candidates, prioritizing those with high lexical overlap with the target examples.
DSIR (Xie et al., 2023)
Data Selection via Importance Resampling (DSIR) is an efficient method that weights candidate examples based on $n$-gram feature overlap with the target validation distribution. A coreset is then formed by resampling from the candidate pool according to these importance weights.
CLD (Nagaraj et al., 2025)
Correlation of Loss Differences (CLD) is a forward-only method that identifies impactful training data by measuring the alignment between a candidate’s training loss trajectory and that of a held-out validation set. The score for each sample is the Pearson correlation between its vector of epoch-to-epoch loss differences and the average vector for the validation set, requiring only per-sample loss values from training checkpoints.
S2L (Yang et al., 2024)
Small-to-Large (S2L) is a scalable data selection method that leverages a small proxy model to guide selection for a larger target model. It first collects the training loss trajectories of all candidate examples by training the small model. It then clusters these trajectories and performs balanced, uniform sampling from the resulting clusters to form the final coreset.
RDS (Hanawa et al., 2021)
Representation-based Data Selection (RDS) uses the model’s hidden representations as features to score data. In our experiments, we follow the implementation in Xia et al. (2024), which uses the final-layer hidden state of the last token in a sequence as its representation. Candidate examples are then scored based on the cosine similarity of their representation to the average representation of the target validation set.
LESS (Xia et al., 2024)
Low-rank gradient Similarity Search (LESS) is an optimizer-aware method for targeted data selection. It adapts the classic gradient-similarity influence formulation to work with the Adam optimizer and variable-length instruction data. To remain efficient, LESS uses LoRA and random projections to compute a low-dimensional "gradient datastore" from a short warmup training phase. Candidates are scored by the cosine similarity of their gradient features to those of the target validation set, aggregated across several checkpoints.
TAGCOS (Zhang et al., 2025b)
Task-Agnostic Gradient Clustered Coreset Selection (TAGCOS) is an unsupervised method that uses sample gradients from a warmed-up model as data representations. The method follows a three-stage pipeline: (1) it computes low-dimensional gradient features for each candidate sample; (2) it performs K-means clustering on these gradient features to group similar data; and (3) it applies an efficient greedy algorithm, Orthogonal Matching Pursuit (OMP), to select a representative subset from within each cluster.
Appendix D Training Datasets Overview
Training data.
Our selection pool $S$ is the union of four public instruction-tuning sources, Dolly, CoT, Oasst1, and Flan_v2, totaling 270,679 examples after filtering (Table 5). We obtain the processed datasets from LESS repository (Xia et al., 2024). Unless otherwise stated, we do not rebalance per-source proportions prior to selection; methods operate over the pooled corpus $S$ with fixed preprocessing and tokenization settings across all experiments.
Dataset #Samples Brief Purpose
Dolly 15,01115{,}011
Instruction following;
everyday tasks, Q&A,
and summaries.
CoT 100,000100{,}000
Chain-of-thought style
prompts and responses.
Oasst1 55,66855{,}668
Multi-turn chat-style
conversations.
Flan_v2 100,000100{,}000
Mixture of diverse tasks
for instruction finetuning.
[…]
Table 5: Source training datasets in the selection pool $S$. “#Samples” reflects post-filter counts.
Appendix E Target Datasets & Evaluation Setup
Evaluation protocol.
We evaluate on four standard targets, CommonsenseQA, SocialIQA, HellaSwag, and Gsm8K, following the OpenAI-style instruction-tuning evaluation protocol (Wang et al., 2023). For selection, we use a small few-shot validation set $T_{\text{val}}$ of five labeled examples per task (ten for Gsm8K), with no access to test labels. At test time, we use fixed in-context exemplars with the shot counts in Table˜6 (5/0/5/8 for CommonsenseQA/SocialIQA/HellaSwag/Gsm8K), sampled from development pools disjoint from $T_{\text{val}}$. We report accuracy for the multiple-choice tasks and exact match on the final numeric answer for Gsm8K. All prompts and decoding settings are held fixed across methods; no task-specific hyperparameters are tuned on test sets. Datasets are obtained via HuggingFace datasets (Wolf et al., 2020), and evaluation is performed with the EleutherAI lm-evaluation-harness (Sutawika et al., 2023).
Task #Val #Test Eval Shots
CommonsenseQA 5 1,221 5
SocialIQA 5 1,954 0
HellaSwag 5 10,042 5
Gsm8K 10 1,319 8
Table 6: Target datasets and evaluation protocol. “#Val” is the few-shot target validation set $T_{\text{val}}$ used for selection; “Eval Shots” is the number of in-context examples at test time.
Appendix F Training Details and Hyperparameters
F.1 Artifact Licenses
According to their license, all the LLMs used in this paper fall under acceptable use cases. The licenses for the models are linked for perusal: Llama-2-7b, Llama-3.2-1B, Llama-3.1-8b and, Mistral 7B (v0.3).
All software dependencies, including PyTorch and torchvision, are open-source and distributed under MIT or BSD-compatible licenses.
F.2 Implementation Details and Shared Hyperparameters
All models are trained with the same setup, closely following LESS (Xia et al., 2024), using an identical training recipe for data loading, preprocessing, optimizer/schedule, precision, and LoRA configuration; only the underlying architecture varies unless otherwise noted.
Optimizer / LR AdamW / $2\times 10^{-5}$
Schedule / Warmup linear / $0.03$
Epochs 4
Per-device BS / Accum 2 / 32 (eff. 64)
Max sequence length 2048
Precision bf16
LoRA (targets) r=128, $\alpha=512$, $p=0.1$ (q,k,v,o_proj)
Eval / Saves no eval; save/1055, keep 15
Table 7: Shared hyperparameters used across all runs; only the architecture varies.
F.3 Model Size and Computational Budget
We evaluate models from 1B to 8B parameters using a shared training/adaptation recipe (Appendix F.2). Concretely, our backbones are Llama-3.2-1B ($\approx$ 1B parameters), Llama-2-7B ($\approx$ 7B parameters), Mistral-7B ($\approx$ 7B parameters), and Llama-3.1-8B ($\approx$ 8B parameters). All runs were executed on a single NVIDIA H200 GPU on an internal cluster (CUDA 12.1, PyTorch 2.1, HuggingFace Transformers 4.46). We do not report wall-clock time or GPU-hours, as these depend on dataset composition and cluster load; instead, we provide asymptotic computational cost and scaling discussion in Section˜5.
Appendix G Ablations on CommonsenseQA (Llama-3.2-1B)
All ablations are run on CommonsenseQA with Llama-3.2-1B at a fixed budget $p=5\%$ and identical fine-tuning hyperparameters. Row entries modify only one factor relative to the default (ours) configuration, which is marked with $\dagger$.
Setting CSQA Acc. (%)
[…]
None 36.27 (0.2)
[…]
0.5 39.35 (0.84)
[…]
Max only 38.54 (0.4)
Mean only 39.02 (0.7)
[…]
Pretrained (no warmup) 37.85 (0.5)
Checkpoint #1 (early) 38.78 (0.3)
[…]
Last 3 38.35 (0.6)
[…]
Last 12 40.68 (0.3)
Table 8: Ablations on the different hyperparameters that can be used for TRIM for choosing a $5\%$ coreset of CommonsenseQA (CSQA) on a Llama-3.2-1B model. Each section tweaks one factor; and all others stay at the default (marked $\dagger$): OOV penalty =1.0=1.0, pooling =0.5=0.5 mean ++ $0.5$ max, coverage $\eta=0.05$, scoring checkpoint == last, layers == last 6. The experimental results reported in the main paper, use these default values.
Appendix H Subset Distribution of Selected Examples
We present the distribution of selected examples across different data selection methods in Figure 5. The three approaches exhibit substantially different preferences for data from each training subset. However, we observe that there is not a monotonic relationship between the proportion of data selected from a subset and overall method performance. This suggests that each subset contains valuable examples for the target tasks, and the primary challenge lies in identifying the most task-relevant instances.
Notably, TAGCOS, being a task-agnostic method, maintains nearly identical selection distributions across all four tasks (approximately 38.5% CoT, 35.0% Flan_v2, 20.7% Oasst1, and 5.8% Dolly). In contrast, both TRIM and LESS demonstrate task-adaptive behavior, substantially varying their selection patterns based on the specific characteristics of each target task.
In our specific observations, we find that TRIM predominantly selects data from the Flan_v2 dataset for both SocialIQA (81.3%) and CommonsenseQA (71.8%). This preference appears intuitive, as both tasks involve question answering with multiple choice or short answer formats, which are abundantly represented in Flan_v2’s diverse instruction-following examples. For GSM8K, TRIM exhibits a strong shift toward the CoT dataset (67.2%), while reducing Flan_v2 to 32.0%. This selection pattern aligns well with GSM8K’s focus on mathematical reasoning, where step-by-step problem solving, a hallmark of the CoT dataset, is essential. Finally, for Hellaswag, TRIM adopts a more balanced approach, selecting comparable amounts from Flan_v2 (56.0%) and CoT (42.7%). This distribution reflects Hellaswag’s nature as a commonsense inference task that benefits from both natural language understanding examples and reasoning-based training data.
Figure 5: Distribution of selected examples across training subsets for different methods and target tasks. Each bar shows the percentage breakdown of data selected from CoT, Dolly, Flan_v2, and Oasst1 datasets.